# TIM: An Efficient Temporal Interaction Module for Spiking Transformer

### v1

@@ 7 Conclusion
In our research, we conduct a detailed examination of the current Spiking Transformer models, focusing particularly on their suboptimal use of temporal information. To address this issue, we introduce the Temporal Interaction Module (TIM), a groundbreaking, plug-and-play component designed to enhance the Spiking Transformer’s ability to process temporal data more effectively. This module, skillfully constructed using one-dimensional convolutions, is notable for its minimal parameter count. Our methodology ensures simplicity in implementation while adhering to the Spiking Transformer’s principles of efficiency and compactness. Rigorous experimental validations underscore the efficacy of TIM, revealing its superior performance, especially prominent in neuromorphic dataset applications. TIM not only excels in these contexts but also establishes new state-of-the-art benchmarks. Furthermore, our comprehensive ablation studies corroborate the effectiveness of TIM in accurately capturing and utilizing temporal information, solidifying its role as a pivotal advancement in neuromorphic computing.

### v3

@@ 7 Conclusion
In our research, we conduct a detailed examination of the current Spiking Transformer models, focusing particularly on their suboptimal use of temporal information. To address this issue, we introduce the Temporal Interaction Module (TIM), a groundbreaking, plug-and-play component designed to enhance the Spiking Transformer’s ability to process temporal data more effectively. This module, skillfully constructed using one-dimensional convolutions, is notable for its minimal parameter count. Our methodology ensures simplicity in implementation while adhering to the Spiking Transformer’s principles of efficiency and compactness. Rigorous experimental validations underscore the efficacy of TIM, revealing its superior performance, especially prominent in neuromorphic dataset applications. TIM not only excels in these contexts but also establishes new state-of-the-art benchmarks. Furthermore, Our discussion and ablation study demonstrate that TIM effectively extracts temporal information on complex datasets, leading to improvements in model performance. Meanwhile, TIM retains the efficiency characteristic of SNNs and the ability to generalize across various Spiking Transformer architectures.
Acknowledgments
This research is financially supported by a funding from Institute of Automation, Chinese Academy of Sciences (Grant No. E411230101).